Having issues with this question please advise.
The scores on a national test were normally distributed with a mean of 83 and a standard deviation of 8. If 400 test takers were selected at random, how many of them would you
expect to have a score lower than 75? (Round to the nearest person)

Answer by stanbon(75887)

The scores on a national test were normally distributed with a mean of 83 and a standard deviation of 8. If 400 test takers were selected at random, how many of them would you
expect to have a score lower than 75? (Round to the nearest person)
-----
z(75) = (75-83)/[8/sqrt(400)] = -8/[8/20] = -20
-----
P(x-bar < 75) = P(z < -20) is practically zero.
==================
Cheers,
Stan H.